an isolated charged point partivle produces elwctric foreld E at 2m away claculate E at 2m away?
1
Expert's answer
2020-11-20T09:35:12-0500

Single point-like charge produce electric field E:

"\displaystyle E = \frac{1}{4 \pi \epsilon_0} \frac{q}{r^2}"

"\displaystyle \frac{1}{4 \pi \epsilon_0} = 8,987 \cdot 10^9 \; [\frac{N \cdot m^2}{C^2}], r=2\, m."

So, depending on the value of charge, E will be

"\displaystyle E= \frac{8.987 \cdot 10^9}{4} q =2.247 q \;[\frac{V}{m}]"